John is currently 7 years older than paul. in two years time john will be twice pauls age. construct a suitable table and equation to find johns age.Can you help me construct both the Table and the equation?
Answer by rothauserc(4718)

let paul's age be x, then we have the following
then john's age is x+7, now in 2 years
(x+7)+2 = 2*(x+2)
x+9 = 2x+4
x = 5
paul's age is 5 and john's age is 12